if cathode A has increased in mass by 1.25g, calculate the mass which would have been deposited by one faraday of electricity at A
1
Expert's answer
2017-06-14T08:16:10-0400
The Faraday’s law of electrolysis:
m=Q/F∙M/z
where m is mass, Q is total electricity passed through the solution, M – molar mass and z – number of ions.
Let’s write equation for the process when cathode A increases:
1.25 g=Q_1/F∙M/z
When 1F passes through solution:
m_2=1/F∙M/z

For both of processes M, F and z remain the same. If we divide first equation by second we will get:
(1.25 g)/m_2 =Q_1/1
The mass of cathode A in second case would be:
m_2=(1.25 g )/Q_1